OpenStudy (anonymous):

Convert -498degrees to radian measure in terms of pie.

OpenStudy (jdoe0001):

pie? sounds delicious! I do know only \(\boldsymbol{\pi}\)i though

OpenStudy (jdoe0001):

\(\bf radians=\cfrac{degrees\cdot \pi}{180}\)

OpenStudy (isaiah.feynman):

Hey spiderman's girl firend, its spelt as "pi" actually. :)

OpenStudy (anonymous):

haha sorry I didn't write it correctly.

OpenStudy (isaiah.feynman):